OpenStudy (anonymous):

Ramanujan discovered 1729 as a magic number. Why 1729 is a magic number ? Brain teaser

OpenStudy (anonymous):

ooh ooh i know

OpenStudy (anonymous):

well actually i forget something like can be written as a sum of squares in two different ways or some such

sammixboo (sammixboo):

I don't. Never knew about this. Would love to hear about it! http://en.wikipedia.org/wiki/1729_%28number%29

OpenStudy (wolf1728):

It is the smallest positive number that can be expressed as the sum of 2 cubes in 2 different ways. 1729= 1,728 + 1 (or 12 cubed + 1 cubed) 1729 = 1,000 + 729 (or 10 cubed + 9cubed)

OpenStudy (anonymous):

lol i was close... squares, cubes, whatever

OpenStudy (anonymous):

Two different set of numbers @satellite73 do you know the numbers

OpenStudy (wolf1728):

lol satellite

sammixboo (sammixboo):

Hehe, so same yet so different, Satellite

OpenStudy (anonymous):

yeah you can't square a circle but you can cube one

sammixboo (sammixboo):

Um 1729 = 1^3 + 12^3 = 9^3 + 10^3 (idk wikipedia)

OpenStudy (alexandervonhumboldt2):

it is magic because: It is 1729. Discovered by mathemagician Srinivas Ramanujan, 1729 is said to be the magic number because it is the sole number which can be expressed as the sum of the cubes of two different sets of numbers.

OpenStudy (alexandervonhumboldt2):

here you go!!!!!!!!!!!!!!!!!

OpenStudy (alexandervonhumboldt2):

@bingtamekia

OpenStudy (anonymous):

what's up @AlexandervonHumboldt2